# Nonlinear Model Order Reduction of Dynamical Systems in Process Engineering: Review and Comparison

### v1

@@ 5.2.  Data sampling
In order to guarantee a fair comparison of snapshot-based methods against each other, we always use the same training data. This data set is recorded by simulating the FOM subject to randomly ordered inputs drawn from an equidistant $5\times 5$ grid in $\mathcal{U}$. Following [240], the sampling time is $\Delta t_{\mathrm{s}}=15\,\mathrm{min}$. We impose the input sequence as steps (zero-order hold) of 240 h (10 days) duration. Thereby, we include information from both the fast spectrum of the system response and the stationary states. We illustrate the input profile $\bm{u}(t)$ in the supplementary information (SI).
[…]
@@ 5.3.  Implementation
As previously discussed, POD and MFL are strongly affected by scaling. Herein, we scale all states and inputs individually to the range $[-1,1]$ and apply zero-mean centering. We remark that a tailored scaling strategy, e.g., scaling based on physical relevance and range, may enable more accurate ROMs (with respect to a target application). Like the FOM, we implement all intrusive ROMs in Modelica and perform all simulations in Dymola using DASSL with an integration tolerance of $10^{-6}$. The non-intrusive ROMs (DMDc, KW, MFL-ANN) are directly evaluated in the respective ML environments in Python (see below). Next, we provide further details on the implementation.
@@ 5.3.1.  POD methods and DMDc
We implement POD and DMDc using SciPy as a Python interface to the SVD implementation in LAPACK. All other matrix computations are performed using Numpy. We specify float64 precision, which was found to be crucial for obtaining reliable SVD results. We scale the latent states $\bm{z}(t)$ heuristically, which was found to improve model conditioning. For DMDc, we calculate the coefficient matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{U}_{1}$ in Eq. 24 using the formulas given in [124]. To derive a POD-Res model, we use a truncated residualization, where $n_{q}=50-n_{z}$. The truncation improves simulation convergence without significant loss of precision.
@@ 5.3.2.  MFL-Galerkin and MFLu-Galerkin methods
We train AENs in Tensorflow using the standard training problem with MSE reconstruction loss [102]. To reduce the number of trainable parameters and speed up the training, we prereduce the scaled and zero-centered snapshots by projecting the data onto the first $100$ POD modes [101]. For the success of MFL, it is crucial to not rescale the projected snapshots again. The encoders and decoders have a single fully connected nonlinear hidden layer with 50 neurons with (smooth) tanh activation, respectively. Using deeper networks or additional neurons did not improve the reconstruction accuracy considerably.
We divide the pre-processed snapshots into 80 % training data and 20 % validation data. For every network, we run three trainings with random seeds and select the weights with the best performance on the validation data set. We train all networks for a total of 20 0002000020\,00020 000 epochs and minibatch size 256, using Adam with initial learning rate (LR) of $10^{-4}$ for the first 15 0001500015\,00015 000 epochs and SGD with LR of $10^{-6}$ for subsequent finetuning. For the implementation of MFLu, we modify the AEN structure from $(\bm{\varphi},\bm{\varphi}^{\dagger})$ to $(\bm{\psi},\bm{\psi}^{\dagger})$ as discussed in Section 3.1.5. However, the general training setup remains the same. To include the AENs in the ROMs, we implement the decoder network structure in Modelica.
@@ 5.3.3.  KW models
We use our ML framework [51] to train reduced KW models. This framework is based on Tensorflow and implements the KW model structure along with a training procedure based on snapshot data. As in Section 5.3.2, we pre-reduce the training data through projection onto the principal $100$ POD modes. The trajectory snippets required for the prediction loss are created by sliding along the training data set in a moving horizon fashion, stopping every 10 samples and copying a series of 100 consecutive snapshots. The encoder and decoder networks have a symmetric structure with two hidden layers and linear output layer. The number of hidden neurons are determined via linear interpolation between $100$ and $n_{z}$. We use ELU activation and train the model for 10 0001000010\,00010 000 epochs and minibatch size 32, where the first 8 00080008\,0008 000 epochs use ADAM with LR of $10^{-3}$ and the final 2 00020002\,0002 000 epochs use SGD with LR of $10^{-6}$ for finetuning. All other hyperparameters have the default values from [51]. The ROM simulation employs the graph-based execution mode of Tensorflow.
@@ 5.3.4.  MFL-ANN models
We implement the method in [159] by means of a modification of our aforementioned framework [51]. To this end, we substitute the linear dynamics in Eq. 26 by a feedforward ANN to learn the discrete-time latent predictor:
@@
where $\bm{F}:\mathcal{Z}\times\mathcal{U}\rightarrow\mathcal{Z}$ and $k\in\{0,1,...\}$. The ANN $\bm{F}$ features two hidden layers of each $2n_{z}$ neurons and ReLU activation, and a linear output layer. This specification is parsimonious compared to [159], who use three hidden layers of each $5n_{z}$ neurons. However, we found that overparameterization results in poor training convergence. All other settings, including AEN structure and training hyperparameters, are identical to Section 5.3.3.
@@ 5.3.5.  COMP and AGG models
Both methods are applied to the ASU process model in a unit-by-unit fashion. However, the submodels of HX3, HPC feed stage, and IRC already have a minimal form and cannot be further reduced by the COMP or AGG. Since the dynamic relevance of these units is unclear without further investigation, we leave them untouched. Moreover, preliminary experimentation showed that the HX dynamics have a noticeable effect on the overall system response. Hence, we do not replace these units by quasi-stationary models, i.e., we create and insert ROM submodels for HX1, HX2, and HPC.
The number and size of compartments and the position of the sensitive elements (referring to HPC trays and HX cells simultaneously) are degrees of freedom of the COMP method and may be regarded as hyperparameters. Similarly, the hyperparameters of AGG are the number and positions of aggregation elements. We follow [183] and use compartments of a uniform size, with sensitive elements located in the center of the respective compartments. A perturbation analysis of the positions confirmed this choice. We select the same central positions for the aggregation elements. Using the same number of compartments or aggregation elements $n_{\mathrm{c}}$ for each HX1, HX2, and HPC, we obtain $n_{z}=10+6\cdot n_{\mathrm{c}}$. In all AGG models, we specify $\mathbf{H}$ in Eq. 27 such that $\mathrm{diag}(\mathbf{H})=n_{x}$.
@@ 5.4.  Definition of the comparison criteria
We compare the selected MOR techniques in two tests. First, we only assess the accuracy of the linear or nonlinear subspaces, i.e., the slow manifold approximations, constructed by the MOR methods. This assessment does not yet include ROMs, but merely determines the average accuracy of projecting the FOM state snapshots of a test data set onto the respective low-dimensional subspaces. In particular, the projection error constitutes a lower bound on the ROM prediction error [22]. Second, we simulate the ROMs, i.e., we examine the prediction capabilities of the ROMs. This prediction test is the most important criterion as we assess the overall ROM accuracy. We initialize all ROMs at the nominal steady state described above.
In both evaluations, we use an independent test data set that comprises a series of random uniformly distributed zero-order hold inputs (illustration see SI). Furthermore, we consider different degrees of reduction, i.e., we vary the reduced order $n_{z}$. As previously discussed, for COMP and AGG, the values of $n_{z}$ are not freely selectable but are restricted by the subsystem-by-subsystem application of the methods, where viable options are $n_{z}\in\{16,22,28,...\}$. To facilitate a comparison of the MOR methods, we include such values of $n_{z}$ into the study.
@@
Therein, $\mathbf{X}\in\mathbb{R}^{n_{x}\times N}$ is the scaled test snapshot matrix, having the snapshots $\bm{x}^{(k)}$, $k=1,...,N$, as matrix rows. The matrix $\mathbf{X}_{\mathrm{r}}\in\mathbb{R}^{n_{x}\times N}$ collects the projected or predicted snapshots, $\bm{x}_{\mathrm{r}}^{(k)}\in\mathbb{R}^{n_{x}}$, $k=1,...,N$. We scale and shift $\mathbf{X}$ and $\mathbf{X}_{\mathrm{r}}$ consistently with the training data.
We note, however, that the choice of the error norm and test data set can have a significant impact on the conclusions and should always be selected based on the ROM application. For example, the RMSE averages out peak deviations so that the slow and steady-state errors are prioritized over maximum errors, e.g., an IR on a short-term scale does not drastically increase the error. Herein, our test data set comprises a significant share of slow response and near-stationary snapshots. We consider this choice to be suitable for examining ROMs that capture the slow manifold.
Finally, we describe how the projections for test 1 are obtained. Regarding the linear basis methods POD and DMDc, both the linear subspaces and the projections coincide, respectively. The projection is given by the projection matrix, $\mathbf{\Pi}\triangleq\mathbf{U}_{1}\mathbf{U}_{1}^{\mathrm{T}}$, where $\mathbf{U}_{1}$ is from Eq. 10 or Eq. 24. To compute the projections performed by MFL, MFLu, KW, and MFL-ANN, we encode and decode the test data through the respective AENs. The projections in POD-Res, COMP and AGG are inherently defined through the degeneration/residualization. We evaluate these projections by solving the respective nonlinear systems of equations, Eqs. 16b, 13b and 27b, on the test data in Dymola. Recall that for the selected hyperparameters, the slow manifolds of AGG and COMP coincide; however, the projections differ.
@@ 5.5.  Results: Projection accuracy
We present the results of the projection test. Table 2 compares the projection errors (RMSE) on the test data set. First, we consider small $n_{z}\leq 10$, i.e., the very-low-order region. As expected, the non-intrusive nonlinear subspace methods (POD-Res, MFL, MFLu) perform significantly better than the linear basis methods (POD, DMDc). The respective difference in RMSE is approximately one order of magnitude. Overall, the smallest projection errors are reached by the MFLu approach. However, the difference in RMSE between MFL and MFLu decreases with increasing $n_{z}$. This confirms the limitations of MFL discussed in Section 3.1.5. Clearly, the error difference is system-dependent and may be even more substantial for a strongly parameterized slow manifold. Finally, the projection RMSE of the non-intrusive nonlinear subspace methods KW and MFL-ANN is comparable to POD. The higher RMSE than for MFL is attributed to the additional loss terms included in the KW and MFL-ANN trainings.
@@
2 $7.8\cdot 10^{-2}$ $4.9\cdot 10^{-2}$ $1.0\cdot 10^{-2}$ $5.3\cdot 10^{-3}$ $1.5\cdot 10^{-1}$ $1.6\cdot 10^{-1}$- --
5 $1.6\cdot 10^{-2}$ $1.7\cdot 10^{-2}$ $2.2\cdot 10^{-3}$ $1.8\cdot 10^{-3}$ $1.4\cdot 10^{-2}$ $2.5\cdot 10^{-2}$- --
10 $3.8\cdot 10^{-3}$ $1.2\cdot 10^{-2}$ $1.3\cdot 10^{-3}$ $1.4\cdot 10^{-3}$ $7.3\cdot 10^{-3}$ $2.5\cdot 10^{-2}$- --
@@
Moving to higher orders, i.e., $n_{z}\in\{16,28\}$, the improvement by the MFL-based methods is only moderate, which was also observed in [36, 101]. We attribute this limited improvement to the challenging convergence of the respective unsupervised learning problems using local solvers, i.e., training AENs by stochastic gradient descent algorithms. As a result, POD truncation outperforms the MFL methods at $n_{z}=28$, facilitating the overall smallest projection error. Interestingly, the linear POD subspace projection also exceeds the accuracy of the nonlinear POD-Res method. A further investigation of the POD-Res results revealed moderate deviations between $\bm{x}(t)$ and $\bm{x}_{\mathrm{r}}(t)$ at the points of abrupt changes in $\bm{u}(t)$. Finally, both COMP and AGG exhibit relatively large errors compared to the other ROM methods. Here, the COMP subspace shows significant deviations from the FOM samples near the input steps but relatively small errors otherwise. In contrast, the maximum deviations of AGG are lower, but a greater number of samples have a notable projection error.
@@ 5.6.  Results: ROM predictions
We now present the results of the ROM prediction test. Table 3 collects the state prediction error (RMSE) of the ROMs. The RMSE is consistently higher than in Table 2, because ROM predictions are less accurate than the (orthogonal) projection of FOM snapshots onto the ROM subspace.
Both linear subspace methods, POD-Galerkin and DMDc, do not reach the desired accuracy at very low order, i.e., $n_{z}\leq 10$. For reduced orders $11\leq n_{z}\leq 21$, the simulations of the POD-Galerkin ROMs did not converge at all. As we discuss in the SI, these critical values of $n_{z}$ lie below the truncation threshold of $n_{z}=23$, obtained from the POD singular values. At order $n_{z}=28$, the POD-Galerkin ROM successfully provides an accurate prediction, whereas the DMDc prediction is still poor. Clearly, the nonlinearity of the FOM prohibits highly accurate DMDc models. Finally, we notice a significant gap between projection error (Table 2) and projection error (Table 3) for both POD-Galerkin and DMDc.
Compared to the POD-Galerkin results, the POD-Res model is more accurate throughout all orders. This advantage is particularly significant in the very-low-order region. In contrast to the projection accuracy (Table 2), POD-Res outperforms the POD-Galerkin ROMs in the prediction. Even for $n_{z}=2$, we notice a fairly small prediction error of POD-Res. Since an RMSE of $10^{-2}$ roughly corresponds to 1 % average error in the scaled states, we expect a satisfactory performance of POD-Res ROMs of order $n_{z}=10$ (and possibly even $n_{z}=5$) in many applications. When moving to higher $n_{z}$, the accuracy improves further.
Table 3: Results of ROM predictions. The RMSE indicates the average error between the state predictions by the ROMs versus the true FOM states. We consider different ROM orders $n_{z}$. Non-converged simulations are indicated by ()(*)( ∗ ).
@@
2 $1.1\cdot 10^{-1}$ $1.5\cdot 10^{-1}$ $3.6\cdot 10^{-2}$ $6.1\cdot 10^{-1}$ $2.9\cdot 10^{-1}$ $1.5\cdot 10^{-1}$ $1.7\cdot 10^{-1}$- --
5 $6.3\cdot 10^{-2}$ $1.5\cdot 10^{-1}$ $2.1\cdot 10^{-2}$ $1.4\cdot 10^{-1}$ $2.2\cdot 10^{-2}$ $1.5\cdot 10^{-2}$ $2.9\cdot 10^{-2}$- --
10 $4.9\cdot 10^{-2}$ $1.3\cdot 10^{-1}$ $8.8\cdot 10^{-3}$ $2.1\cdot 10^{-2}$ $1.5\cdot 10^{-2}$ $9.9\cdot 10^{-3}$ $3.3\cdot 10^{-2}$- --
16 ()(*)( ∗ ) $9.4\cdot 10^{-2}$ $7.9\cdot 10^{-3}$ $1.0\cdot 10^{-2}$ $1.1\cdot 10^{-2}$ $1.9\cdot 10^{-2}$ $3.2\cdot 10^{-2}$ $2.5\cdot 10^{-1}$ $4.9\cdot 10^{-2}$
28 $8.7\cdot 10^{-3}$ $7.8\cdot 10^{-2}$ $7.5\cdot 10^{-3}$ ()(*)( ∗ ) ()(*)( ∗ ) $1.1\cdot 10^{-2}$ $3.3\cdot 10^{-2}$ $3.7\cdot 10^{-2}$ $1.7\cdot 10^{-2}$
@@
The non-intrusive KW and MFL-ANN ROMs have a prediction error comparable to the intrusive nonlinear subspace methods. The differences in RMSE lie within the range of deviations found in ML trainings. Notably, the MFL-ANN models do not perform better than the KW models. A comparison to the performance on the training data set revealed a tendency of MFL-ANN to overfit the training data (the RMSEs on the training data were approximately one order of magnitude smaller than on the test data), which was less pronounced for the KW structure. From $n_{z}=10$ to $n_{z}=28$, there is no improvement in RMSE for both model types, i.e., increasing the ROM order does not improve the accuracy. Finally, we notice that the projection and prediction errors are comparable for both models, respectively.
For COMP and AGG models with a single compartment and aggregation cell for each unit operation, respectively, we find significant prediction errors. In particular, the COMP model is far from the desired accuracy due to severe IR. Increasing the ROM order results in a notably smaller RMSE, where the AGG results reach the desired error level. However, the COMP response still exhibits a notable IR, diminishing ROM accuracy (see figure below).
@@
Fig. 5: Results of test 2 (prediction). Depicted is the liquid molar fraction of oxygen on the HPC top stage over the first 2000 h of the test. ROM order is (a) $n_{z}=5$ and (b) $n_{z}=28$.
To illustrate the results, we examine the molar fraction of oxygen on the HPC top tray, which is directly connected to the quality of the nitrogen product. We show the predictions for $n_{z}=5$ in Fig. 5(a) and for $n_{z}=28$ in Fig. 5(b). In Fig. 5(a), we observe a significant offset between ROM prediction and FOM response for POD, DMDc, and MFL. This finding is consistent with Table 3. For these ROMs, the offset involves both a transient and a steady-state mismatch. Moreover, despite the relatively small RMSE of POD-Res, the graph shows significant IR near the input steps, whereas the steady states are captured accurately. All other ROMs (MFLu, KW, MFL-ANN) reproduce the FOM trajectory closely.
Observing the higher-order case in Fig. 5(b), we find significantly better predictions by the POD and POD-Res models. At the same time, the DMDc forecast remains inaccurate. The effect of overfitting discussed above for the MFL-ANN ROMs is also visible in the figure. Compared to Fig. 5(a), the prediction is less accurate despite a higher-order model. Further, notable deviations from the FOM trajectory are visible for the COMP model in the form of severe IR. In contrast, the AGG ROM provides a close reproduction of the FOM trajectory.

### v2

@@ 5.2. Data sampling
In order to guarantee a fair comparison of snapshot-based methods against each other, we always use the same training data. This data set is recorded by simulating the FOM subject to randomly ordered inputs drawn from an equidistant $5\times 5$ grid in $\mathcal{U}$. Following [44], the sampling time is $\Delta t_{\mathrm{s}}=15\,\mathrm{min}$. We impose the input sequence as steps (zero-order hold) of 240 h (10 days) each. Thereby, we include information from both the fast spectrum of the system response and the stationary states. We illustrate the input profile $\bm{u}(t)$ in the supplementary information.
We simulate the FOM in Dymola 2025x using the integrator DASSL [224] with integration tolerance $10^{-6}$. The recorded data set comprises N=24  [^0] N=24\,000 snapshots of each state. The corresponding snapshot matrices have the form $\mathbf{X}\in\mathbb{R}^{n_{x}\times N}$ and $\mathbf{U}\in\mathbb{R}^{n_{u}\times N}$. These snapshots cover the operating space between 15 mol/s and 27 mol/s production rate and product quality between 10 ppm and 1000 ppm impurities.
@@ 5.3. Implementation
As previously discussed, POD and MFL are strongly affected by scaling. Herein, we scale all states and inputs individually to the range [1,1][-1,1] and apply zero-mean centering. We remark that a tailored scaling strategy, e.g., scaling based on physical relevance and range, may enable more accurate ROMs (with respect to a target application). Like the FOM, we implement all intrusive ROMs in Modelica and perform all simulations in Dymola using DASSL with an integration tolerance of $10^{-6}$. The non-intrusive ROMs (DMDc, KW, MFL-ANN) are directly evaluated in the respective machine learning environments in Python (see below). Next, we provide further details on the implementation.
@@ 5.3.1. POD methods and DMDc
We implement POD and DMDc using SciPy as a Python interface to the SVD implementation in LAPACK. All other matrix computations are performed using Numpy. We specify float64 precision, which was found to be crucial for obtaining reliable SVD results. We scale the latent states $\bm{z}(t)$ heuristically, which was found to improve model conditioning. For DMDc, we calculate the coefficient matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{U}_{1}$ in (25) using the formulas given in [229]. To derive a POD-Res model, we use a truncated residualization, where $n_{q}=50-n_{z}$. The truncation improves simulation convergence without significant loss of precision.
@@ 5.3.2. MFL-Galerkin and MFLu-Galerkin methods
We train AENs in Tensorflow using the standard training problem [286]. To reduce the number of trainable parameters and speed up the training, we prereduce the scaled and zero-centered snapshots by projecting the data onto the first $100$ POD modes [227]. For the success of MFL, it is crucial to not rescale the POD-projected snapshots again. The encoders and decoders have a single fully connected nonlinear hidden layer with 50 neurons with (smooth) tanh activation, respectively. Using deeper networks or additional neurons did not improve the reconstruction accuracy considerably.
We divide the pre-processed snapshots into 80 % training data and 20 % validation data. For every network, we run three trainings with random seeds and select the weights with the best performance on the validation data set. We train all networks for a total of 20 00020\,000 epochs and minibatch size 256, using Adam with initial learning rate of $10^{-4}$ for the first 15 00015\,000 epochs and SGD with learning rate of $10^{-6}$ for subsequent finetuning. For the implementation of MFLu, we modify the AEN structure from $(\bm{\varphi},\bm{\varphi}^{\dagger})$ to $(\bm{\psi},\bm{\psi}^{\dagger})$ as discussed in Section 3.1.5. However, the general training setup remains the same. To include the AENs in the ROMs, we implement the decoder network structure in Modelica.
@@ 5.3.3. KW models
We use our machine learning framework [261] to train reduced KW models. This framework is based on Tensorflow and implements the KW model structure along with a training procedure based on snapshot data. As in Section 5.3.2, we pre-reduce the training data through projection onto the principal $100$ POD modes. The trajectory snippets required for the prediction loss are created by sliding along the training data set in a moving horizon fashion, stopping every 10 samples and copying a series of 100 consecutive snapshots. The encoder and decoder networks have a symmetric structure with two hidden layers and linear output layer. The number of hidden neurons are determined via linear interpolation between $100$ and $n_{z}$. We use ELU activation and train the model for 10 00010\,000 epochs and minibatch size 32, where the first 8 0008\,000 epochs use ADAM with learning rate of $10^{-3}$ and the final 2 0002\,000 epochs use SGD with learning rate of $10^{-6}$ for finetuning. All other hyperparameters have the default values from [261]. The ROM simulation employs the graph-based execution mode of Tensorflow.
@@ 5.3.4. MFL-ANN models
We implement the method in [187] by means of a modification of our aforementioned framework [261]. To this end, we substitute the linear dynamics in (27) by a feedforward ANN to learn the discrete-time latent predictor:
@@
where $\bm{F}:\mathcal{Z}\times\mathcal{U}\rightarrow\mathcal{Z}$ and $k\in\{0,1,$...\}. The ANN $\bm{F}$ features two hidden layers of each $2n_{z}$ neurons and ReLU activation, and a linear output layer. This specification is parsimonious compared to [187], who use three hidden layers of each $5n_{z}$ neurons. However, we found that overparameterization results in poor training convergence. All other settings, including AEN structure and training hyperparameters, are identical to (5.3.3).
@@ 5.3.5. COMP and AGG models
Both methods are applied to the FOM in a unit-by-unit fashion. However, the submodels of HX3, column feed stage, and reboiler/condenser already have a minimal form and cannot be further reduced by the COMP or AGG. Since the dynamic relevance of these units is unclear without further investigation, we leave them untouched. Moreover, preliminary experimentation showed that the HX dynamics have a noticeable effect on the overall system response. Hence, we do not replace these units by quasi-stationary models, i.e., we create and insert ROM submodels for HX1, HX2, and column.
The number and size of compartments and the position of the sensitive elements (referring to column trays and HX cells simultaneously) are degrees of freedom of the COMP method and may be regarded as hyperparameters. Similarly, the hyperparameters of AGG are the number and positions of aggregation elements. We follow [254] and use compartments of a uniform size, with sensitive elements located in the center of the respective compartments. A perturbation analysis of the positions confirmed this choice. We select the same central positions for the aggregation elements. Using the same number of compartments or aggregation elements $n_{\mathrm{c}}$ for each HX1, HX2, and column, we obtain $n_{z}=10+6\cdot n_{\mathrm{c}}$. In all AGG models, we specify $\mathbf{H}$ in (28) such that $\mathrm{diag}(\mathbf{H})=n_{x}$.
@@ 5.4. Definition of the comparison criteria
We compare the selected MOR techniques in two tests. First, we only assess the accuracy of the linear or nonlinear subspaces, i.e., the slow manifold approximations, constructed by the MOR methods. This assessment does not yet include ROMs but merely determines the average accuracy of projecting the FOM state snapshots of a test data set onto the respective low-dimensional subspaces. In particular, the projection error constitutes a lower bound on the ROM prediction error [235]. Second, we simulate the ROMs, i.e., we examine the prediction capabilities of the ROMs. This prediction test is the most important criterion as we assess the overall ROM accuracy. We initialize all ROMs at the nominal steady state described above.
In both evaluations, we use an independent test data set that comprises a series of random uniformly distributed zero-order hold inputs (illustration see supplementary information). Furthermore, we consider different degrees of reduction, i.e., we vary the reduced order $n_{z}$. As previously discussed, for COMP and AGG, the values of $n_{z}$ are not freely selectable but are restricted by the subsystem-by-subsystem application of the methods, where viable options are $n_{z}\in\{16,22,28,$...\}. To facilitate a comparison of the MOR methods, we include such values of $n_{z}$ into the study.
@@
Therein, $\mathbf{X}\in\mathbb{R}^{n_{x}\times N}$ is the scaled test snapshot matrix, having the snapshots $\bm{x}^{(k)}$, k=1,,Nk=1,...,N, as matrix rows. The matrix $\mathbf{X}_{\mathrm{r}}\in\mathbb{R}^{n_{x}\times N}$ collects the projected or predicted snapshots, $\bm{x}_{\mathrm{r}}^{(k)}\in\mathbb{R}^{n_{x}}$, k=1,,Nk=1,...,N. We scale and shift $\mathbf{X}$ and $\mathbf{X}_{\mathrm{r}}$ consistently with the training data.
We note, however, that the choice of the error norm and test data set can have a significant impact on the conclusions and should always be selected based on the ROM application. For example, the RMSE averages out peak deviations so that the slow and steady-state errors are prioritized over maximum errors, e.g., an inverse response on a short-term scale does not drastically increase the error. Herein, our test data set comprises a significant share of slow response and near-stationary snapshots. We consider this choice to be suitable for examining ROMs that capture the slow manifold.
Finally, we describe how the projections for test 1 are obtained. Regarding the linear basis methods POD and DMDc, both the linear subspaces and the projections coincide, respectively. The projection is given by the projection matrix, $\mathbf{\Pi}\triangleq\mathbf{U}_{1}\mathbf{U}_{1}^{\mathrm{T}}$, where $\mathbf{U}_{1}$ is from (11) or (25). To compute the projections performed by MFL, MFLu, KW, and MFL-ANN, we encode and decode the test data through the respective AENs. The projections in POD-Res, COMP and AGG are inherently defined through the degeneration/residualization. We evaluate these projections by solving the respective nonlinear systems of equations, Eqs. (17b), (14b) and (28b), on the test data in Dymola. Recall that for the selected hyperparameters, the slow manifolds of AGG and COMP coincide; however, the projections differ.
@@ 5.5. Results: Projection accuracy
We present the results of the projection test. Table 3 compares the projection RMSE on the test data set. First, we consider small $n_{z}\leq 10$, i.e., the very-low-order region. As expected, the nonlinear subspace methods (POD-Res, MFL, MFLu) perform significantly better than the linear basis methods (POD, DMDc). The respective difference in RMSE is approximately one order of magnitude. Overall, the smallest projection errors are reached by the MFLu approach. However, the difference in RMSE between MFL and MFLu decreases with increasing $n_{z}$. This confirms the input-dependency of the slow manifold as discussed in Section 3.1.5. Clearly, the error difference is system-dependent and may be even more substantial for a strongly parameterized slow manifold. Finally, the projection RMSEs of the non-intrusive nonlinear subspace methods KW and MFL-ANN are comparable to POD. The higher RMSE than for MFL is attributed to the additional loss terms included in the KW and MFL-ANN trainings.
@@
2 $7.8\cdot 10^{-2}$ $4.9\cdot 10^{-2}$ $1.0\cdot 10^{-2}$ $5.3\cdot 10^{-3}$ $1.5\cdot 10^{-1}$ $1.6\cdot 10^{-1}$ - -
5 $1.6\cdot 10^{-2}$ $1.7\cdot 10^{-2}$ $2.2\cdot 10^{-3}$ $1.8\cdot 10^{-3}$ $1.4\cdot 10^{-2}$ $2.5\cdot 10^{-2}$ - -
10 $3.8\cdot 10^{-3}$ $1.2\cdot 10^{-2}$ $1.3\cdot 10^{-3}$ $1.4\cdot 10^{-3}$ $7.3\cdot 10^{-3}$ $2.5\cdot 10^{-2}$ - -
@@
Moving to higher orders, i.e., $n_{z}\in\{16,28\}$, the improvement by the MFL-based methods is only moderate, which was also observed in [159, 227]. We attribute this limited improvement to the challenging convergence of the respective unsupervised learning problems using local solvers, i.e., training AENs by stochastic gradient descent algorithms. As a result, POD truncation outperforms the MFL methods at $n_{z}=28$, facilitating the overall smallest projection error. Interestingly, the linear POD subspace projection also exceeds the accuracy of the nonlinear POD-Res method. A further investigation of the POD-Res results revealed notable deviations between $\bm{x}(t)$ and $\bm{x}_{\mathrm{r}}(t)$ at the points of abrupt changes in $\bm{u}(t)$. Finally, both COMP and AGG exhibit relatively large errors compared to the other ROM methods. Here, the COMP subspace shows significant deviations from the FOM samples near the input steps but relatively small errors otherwise. In contrast, the maximum deviations of AGG are lower, but a greater number of samples have a notable projection error.
@@ 5.6. Results: ROM predictions
We now present the results of the ROM prediction test. Table 4 collects the state prediction RMSE of the ROMs. The RMSE in Table 4 is consistently higher than in Table 3, because ROM predictions are less accurate than the mere projection of FOM snapshots onto the ROM subspace.
Both linear subspace methods, POD-Galerkin and DMDc, do not reach the desired accuracy at very low order, i.e., $n_{z}\leq 10$. For reduced orders $11\leq n_{z}\leq 21$, the simulations of the POD-Galerkin ROMs did not converge at all. As we discuss in the supplementary information, these critical values of $n_{z}$ lie below the truncation threshold of $n_{z}=23$, obtained from the POD singular values. At order $n_{z}=28$, the POD-Galerkin ROM successfully provides an accurate prediction, whereas the DMDc prediction is still poor. Clearly, the nonlinearity of the FOM prohibits highly accurate DMDc models. Finally, we notice a significant gap between projection error (Table 3) and projection error (Table 4) for both POD-Galerkin and DMDc.
Compared to the POD-Galerkin results, the POD-Res model is more accurate throughout all orders. This advantage is particularly significant in the very-low-order region. In contrast to the projection accuracy (Table 3), POD-Res outperforms the POD-Galerkin ROMs in the prediction. Even for $n_{z}=2$, we notice a fairly small prediction error of POD-Res. Since an RMSE of $10^{-2}$ roughly corresponds to 1 % average error in the scaled states, we expect a satisfactory precision of POD-Res ROMs of order $n_{z}=10$ (and possibly even $n_{z}=5$) in tasks such as control. When moving to higher $n_{z}$, the accuracy improves further.
Table 4: Results of ROM predictions. The RMSE indicates the average error between the state predictions by the ROMs versus the true FOM states. We consider different ROM orders $n_{z}$. Non-converged simulations are indicated by ()(*).
@@
2 $1.1\cdot 10^{-1}$ $1.5\cdot 10^{-1}$ $3.6\cdot 10^{-2}$ $6.1\cdot 10^{-1}$ $2.9\cdot 10^{-1}$ $1.5\cdot 10^{-1}$ $1.7\cdot 10^{-1}$ - -
5 $6.3\cdot 10^{-2}$ $1.5\cdot 10^{-1}$ $2.1\cdot 10^{-2}$ $1.4\cdot 10^{-1}$ $2.2\cdot 10^{-2}$ $1.5\cdot 10^{-2}$ $2.9\cdot 10^{-2}$ - -
10 $4.9\cdot 10^{-2}$ $1.3\cdot 10^{-1}$ $8.8\cdot 10^{-3}$ $2.1\cdot 10^{-2}$ $1.5\cdot 10^{-2}$ $9.9\cdot 10^{-3}$ $3.3\cdot 10^{-2}$ - -
16 ()(*) $9.4\cdot 10^{-2}$ $7.9\cdot 10^{-3}$ $0\cdot 10^{-2} 1.1$$1\cdot 10^{-2}$ $1.9\cdot 10^{-2}$ $3.2\cdot 10^{-2}$ $2.5\cdot 10^{-1}$ $4.9\cdot 10^{-2}$
28 $8.7\cdot 10^{-3}$ $7.8\cdot 10^{-2}$ $7.5\cdot 10^{-3}$ ()(*) ()(*) $1.1\cdot 10^{-2}$ $3.3\cdot 10^{-2}$ $3.7\cdot 10^{-2}$ $1.7\cdot 10^{-2}$
@@
The non-intrusive KW and MFL-ANN ROMs have a prediction error comparable to the intrusive nonlinear subspace methods. For each $n_{z}$, the differences in RMSE are insignificant, where KW models perform only slightly better. Notably, the MFL-ANN models do not perform better than the KW models. A comparison to the performance on the training data set revealed a tendency of MFL-ANN to overfit the training data (the RMSEs on the training data were approximately one order of magnitude smaller than on the test data), which was less pronounced for the KW structure. From $n_{z}=10$ to $n_{z}=28$, there is no improvement in RMSE for both model types, i.e., increasing the ROM order does not improve the accuracy. Finally, we notice that the projection and prediction errors are comparable for both models, respectively.
For COMP and AGG models with a single compartment and aggregation cell for each unit operation, respectively, we find significant prediction errors. In particular, the COMP model is far from the desired accuracy due to severe inverse response. Increasing the ROM order results in a notably smaller RMSE, where the AGG results reach the desired error level. However, the COMP response still exhibits a notable inverse response, diminishing ROM accuracy (see Fig. 4 below).
@@
Fig. 4: Results of test 2 (prediction). Depicted is the liquid molar fraction of oxygen on the column top stage over the first 2000 h of the test. ROM order is (a) $n_{z}=5$ and (b) $n_{z}=28$.
To illustrate the results, we examine the molar fraction of oxygen on the column top tray, which is directly connected to the quality of the nitrogen product. We show the predictions for $n_{z}=5$ in Fig. 4(a) and for $n_{z}=28$ in Fig. 4(b). In Fig. 4(a), we observe a significant offset between ROM prediction and FOM response for POD, DMDc, and MFL. This finding is consistent with Table 4. For these ROMs, the offset involves both a transient and a steady-state mismatch. Moreover, despite the relatively small RMSE of POD-Res, the graph shows significant inverse response near the input steps, whereas the steady states are captured accurately. All other ROMs (MFLu, KW, MFL-ANN) reproduce the FOM trajectory closely.
Observing the higher-order case in Fig. 4(b), we find significantly better predictions by the POD and POD-Res models. At the same time, the DMDc forecast remains inaccurate. The effect of overfitting discussed above for the MFL-ANN ROMs is also visible. Compared to Fig. 4(a), the MFL-ANN prediction is less accurate despite a higher-order model. Further, notable deviations from the FOM trajectory are visible for the COMP model in the form of severe inverse response. In contrast, the AGG ROM provides a close reproduction of the FOM trajectory.